OpenStudy (anonymous):

The table below shows the values of f(x) and g(x) for different values of x. One of the functions is a quadratic function and the other is an exponential function. Which function is most likely increasing exponentially?

OpenStudy (anonymous):

x f(x) g(x) 1 3 3 2 6 9 3 11 27 4 18 81 5 27 243 f(x) because it eventually exceeds g(x) g(x) because it eventually exceeds f(x) f(x) because it eventually intersects g(x) g(x) because it will not intersect f(x)

OpenStudy (radar):

f(x) = x^2 + 2 g(x) = 3^x That should answer the question.